# Matching-Based Policy Learning

### v1

@@ 6 Discussion
In this paper, we proposed MB-learning for learning the optimal policy in observational studies, leveraging matching estimators in causal inference literature. Specifically, we developed a (bias-corrected) matching-based estimator of the advantage function and estimated the optimal policy by directly maximizing the resulting estimate over a class of policies. The proposed method possesses robustness in observational studies by circumventing the use of inverse probabilities. Theoretically, we established a finite sample bound for the regret of the learned policy, providing rigorous performance guarantee for MB-learning. Empirically, we demonstrated the competitive performance of the proposed method, particularly when dealing with extreme weights or relatively small sample sizes.
There are several interesting directions this work may be extended. First, MB-learning is intended for low-dimensional settings with binary treatment. In causal inference literature, Antonelli et al. (2018) proposed a doubly robust matching estimator for handling high-dimensional covariates, leveraging variable selection techniques for matching with estimated propensity score and prognostic score. In addition, Wu et al. (2022) developed a matching estimator of average causal exposure-response function based on generalized propensity score. These estimators offer inspiration for adapting MB-learning to different scenarios. Second, we developed MB-learning with a fixed number of matches $M$, and our discussion in Section 3 demonstrated that a larger $M$ yields a lower regret bound. Lin et al. (2023) established a remarkable result that bias-corrected matching estimator (Abadie and Imbens, 2011) with a diverging $M$ exhibits doubly robustness and semiparametrically efficiency under regularity conditions. Hence, developing a framework for MB-learning with a diverging $M$ could yield a robust policy learning approach with an improved regret bound. Finally, the robustness of MB-learning can be attributed to the covariates balancing between the groups, which is achieved through directly matching on covariates. It would be of considerable interest to integrate more sophisticated covariate balancing techniques for learning the optimal policy (e.g., Wong and Chan, 2018; Wang and Zubizarreta, 2020; Fan et al., 2022).
Supplementary Materials
@@ Appendix A Technical Proofs
Before the proofs, we introduce some notations here. Let $(\Theta,d)$ denote a semimetric space. For any $\epsilon>0$, let $N(\epsilon,\Theta,d)$ denote the covering number of $\Theta$, i.e., the minimal number of $d$-balls with radius $\epsilon$ needed to cover $\Theta$. We call a set of points is $\epsilon$-separated if the distance between every pair of points is strictly larger than $\epsilon$, and let $D(\epsilon,\Theta,d)$ denote the packing number of $\Theta$, i.e., the maximum number of points in $\Theta$ such that they are $\epsilon$-separated. The entropy numbers are the logarithms of covering and packing numbers respectively. These definitions can be found elsewhere (e.g., van der Vaart and Wellner, 2023). Throughout the proofs, we use $C$ to denote a generic positive constant, which may change from place to place.
@@ A.1 Proof of Lemma 1
Let $\mathbf{Y}=(Y_{1},\ldots,Y_{n})^{\top}$, note that
@@
where the third equality follows from the fact that, if $M_{il}=1$, then $l\in\mathcal{J}_{M}(i)$ and thus we have $W_{i}=1-W_{l}$ by the definition of $\mathcal{J}_{M}(i)$. Combining these results, we have
@@ A.2 Proof of Lemma 2
Note that $|K_{M}(\pi,i)|\leq\sum_{j=1}^{n}M_{ji}=K_{M}(i)$, it follows directly from the Lemma 3 in Abadie and Imbens (2006). ∎
@@ A.3 Proof of Lemma 3
Let $\mathbf{\hat{Y}}(1)=\left(\hat{Y}_{1}(1),\ldots,\hat{Y}_{n}(1)\right)^{\top}$ and $\mathbf{\hat{Y}}(0)=\left(\hat{Y}_{1}(0),\ldots,\hat{Y}_{n}(0)\right)^{\top}$ denote the $n\times 1$ vectors of matching estimators for potential outcomes under each treatment, then we have
@@
Furthermore, let $\bm{\mu}\mathbf{(X\text{,}W)}=\left(\mu(X_{1},W_{1}),\ldots,\mu(X_{n},W_{n})\right)^{\top}$, we have
@@
Let $\bm{\phi}$ be the $n\times 1$ vector taking value in $\{-1,1\}^{n}$, whose $i$th element is $12\pi(X_{i})-$ 1, and $\bm{\epsilon}$ be the $n\times 1$ with the $i$th element $\epsilon_{i}$ equal to $Y_{i}-\mu(X_{i},W_{i})$. Then we have
@@
where the third equality follows form (6) and $\mathbf{I_{n}}$ is the $n\times n$ identity matrix. Note that the elements of the two $n\times n$ matrices $\mathbf{G_{1}}-\mathbf{I_{n}}$ and $\mathbf{G_{0}}-\mathbf{I_{n}}$ are given by
@@
For each row $i$ such that $W_{i}=0$, we have $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}\bm{\mu}\mathbf{(X\text{,}1)}=\frac{1}{M}\sum_{j\in\mathcal{J}_{M}(i)}\mu(X_{j},1)-\mu(X_{i},1)$, where $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}$ is the $i$th row of $(\mathbf{G_{1}}-\mathbf{I_{n}})$; for each row $i$ such that $W_{i}=1$, we have $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}\bm{\mu}\mathbf{(X\text{,}1)}=0$. Hence, we have
@@
by the similar argument. Furthermore, it follows that
@@
After some algebra and using the similar argument in the proof of Lemma 1, we have
@@ A.4 Proof of Lemma 4
Using the same argument in Abadie and Imbens (2006, Proof of Theorem 1, Page 258), we have $C^{2}n^{2/p}E\left[\frac{1}{n}\sum_{i=1}^{n}\|U_{i,M}\|^{2}\right]=O(1)$. Therefore, we have $B_{M}(\pi)=O_{p}\left(n^{-1/p}\right)$ by Markov’s inequality.
@@ A.5.1 Proof of Theorem 1 (i)
From the result in Lemma 3, we have
@@
We analyze each of the three terms respectively. For the first term, under the Assumption that $\mu(x,w)$ is Lipschitz on a bounded set, then $\sup_{x,w}\mu(x,w)$ is bounded. By week law of large number (van der Vaart, 1998, Proposition 2.16), $\bar{A}(\pi)-A(\pi)\stackrel{{\scriptstyle p}}{{\longrightarrow}}0$. For the second term, let $\bar{\sigma}^{2}=\sup_{x,w}\sigma^{2}(x,w)$, it follows form the Lipschitz assumption that $\bar{\sigma}^{2}$ is bounded. Note that
@@
$\leq$ $E\left[\left(1+\frac{K_{M}(i)}{M}\right)^{2}\sigma^{2}(X_{i},W_{i})\right]=O(1),$
where the second equality follows from that for $i\neq j$, $E\left[\epsilon_{i}\epsilon_{j}|\{X_{i},W_{i}\}_{i=1}^{n}\right]=0$; the inequality follows from triangle inequality and $|K_{M}(\pi,i)|\leq K_{M}(i)$; and the last equality follows from the boundness of $\bar{\sigma}^{2}$ and Lemma 3 in Abadie and Imbens (2006). Hence, we have $E_{M}(\pi)=O_{p}\left(n^{-1/2}\right)=o_{p}(1)$ by Markov’s inequality. For the third term, it follows from Lemma 4 that $B_{M}(\pi)=O_{p}\left(n^{-1/p}\right)=o_{p}(1)$.
@@ A.5.2 Proof of Theorem 1 (ii)
Since
$\sqrt{n}\left(\hat{A}_{match}(\pi)-B_{M}(\pi)-A(\pi)\right)=\sqrt{n}\left(\bar{A}(\pi)-A(\pi)\right)+\sqrt{n}E_{M}(\pi),$
@@
For the second term, we consider the similar argument in the proof of Abadie and Imbens (2006, Theorem 4). Let $E_{M,i}(\pi)=(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i}$. Conditional on $\{X_{i},W_{i}\}_{i=1}^{n}$, $K_{M}(\pi,i)$ is deterministic. Hence, the unit-level terms $\{E_{M,i}(\pi)\}^{n}_{i=1}$ are independently and not identically distributed with mean zero. We proceed to proof the asymptotic normality of $\sqrt{n}E_{M}(\pi)/\sqrt{V^{E}(\pi)}=\sum_{i=1}^{n}E_{M,i}(\pi)\sqrt{nV^{E}(\pi)}$ using Lindeberg-Feller central limit theorem (Billingsley, 1995, Theorem 27.2). Specifically, given $\{X_{i},W_{i}\}_{i=1}^{n}$, we need to verify Lindeberg condition:
@@ A.6 Proof of Corollary 2
We only need to verify that $\sqrt{n}\left(B_{M}(\pi)-\hat{B}_{M}(\pi)\right)=o_{p}(1)$. To do this, we need to use the result of unit-level bias correction (Abadie and Imbens, 2011, Lemma A.2), which is given as follows:
@@ A.7 Proof of Theorem 3
Since $R(\hat{\pi})=\left(A(\pi^{*})-A(\hat{\pi})\right)/2$ by definition, it suffices to analyze $A(\pi^{*})-A(\hat{\pi})$ throughout the proof. Let $\Delta(\pi_{1},\pi_{2})=A(\pi_{1})-A(\pi_{2})$, $\bar{\Delta}(\pi_{1},\pi_{2})=\bar{A}(\pi_{1})-\bar{A}(\pi_{2})$, and $\hat{\Delta}(\pi_{1},\pi_{2})=\hat{A}^{bc}_{match}(\pi_{1})-\hat{A}^{bc}_{match}(\pi_{2})$. It follows from triangle inequality that
@@
We will derive the high probability bounds for the three terms in (A.7) in Step 1-3 separately, and then combine these bounds to complete the proof in Step 4.
@@
To derive this bound, we utilize Lemma 2 in Zhou et al. (2023). First, we rewrite $A(\pi)$ and $\bar{A}(\pi)$ in a vector-based form in order to leverage their Lemma 2. Let
@@
We bound (I) using Hoeffding’s inequality (Vershynin, 2018, Theorem 2.2.6). Under the assumption that $|Y_{i}|\leq L$, we have $|\epsilon_{i}|=|Y_{i}-E[Y_{i}|X_{i},W_{i}]|\leq 2L$ by triangle inequality. In addition, note that $E[(2W_{i}-1)\epsilon_{i}]=E\left[(2W_{i}-1)E\left[\epsilon_{i}\mid X_{i},W_{i}\right]\right]=0$. By Hoeffding’s inequality, for any $t>0$,
@@
Let $\delta=\exp\left(\frac{-t^{2}}{8L^{2}}\right)$, we have
@@
Let $\delta=\exp\left(\frac{-M^{2}t^{2}}{8L^{2}}\right)$, it follows that
@@
Let $\delta=\exp\left(-\frac{t^{2}}{64L^{2}}\right)$, combining (19) and (20) we have
@@
It follows from triangle inequality that
@@
Recall that in (25), we have showed that $G^{0}_{n}(\pi)$ is a sub-Gaussian process given $\{X_{i},W_{i},\epsilon_{i}\}^{n}_{i=1}$. Note that $\|\tilde{\pi}_{k}-\tilde{\pi}_{k-1}\|_{n}\leq\|\tilde{\pi}_{k}-\pi\|_{n}+\|\tilde{\pi}_{k-1}-\pi\|_{n}\leq 3\tilde{D}2^{-k}$, we have
@@
Therefore, let $c_{6}>0$, there exists $N(c_{6},q)$ such that for all $n>N(c_{6},q)$, we have $E\left[\max_{i\leq n}K_{M}(i)\right]\leq c_{6}n^{\frac{1}{q}}$. Combining (24), (29) and (30), for any $q>1$ and $n>N(c_{6},q)$, we have
@@
Conditional on $\{X_{i},W_{i}\}_{i=1}^{n}$, let $b_{i}(\pi)=(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]$, the term
@@
is the sum of independence and $E\left[b_{i}(\pi)\epsilon_{i}\right]=0$. Note that
@@
$V=\underset{\pi\in\Pi}{\sup}\sum_{i=1}^{n}E_{\epsilon}\left[\frac{b^{2}_{i}(\pi)\epsilon^{2}_{i}}{4L^{2}\cdot\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}\right]\leq\underset{\pi\in\Pi}{\sup}\frac{\sum_{i=1}^{n}b^{2}_{i}(\pi)}{\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}\leq\frac{\sum_{i=1}^{n}K_{M}^{2}(i)}{\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}$
@@
2($2\sqrt{\left(2V+16E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]\right)t}\leq$ $2\sqrt{2Vt}+8\sqrt{E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]t}$
@@
Without loss of generality, we let $\alpha=1$. It follows that, conditional on $\{X_{i},W_{i}\}^{n}_{i=1}$, with at least probability $1-\exp(-t)$, we have
@@
Since $\{b_{i}(\pi)\}_{i=1}^{n}$ are constants given $\{X_{i},W_{i}\}^{n}_{i=1}$, we replace the $E_{\epsilon}\left[\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]$ above by $E\left[\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]$. Taking expectation w.r.t. $\{X_{i},W_{i}\}^{n}_{i=1}$, for any $t>0$, with at least probability $1-\exp(-t)$, we have
$\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\leq 2E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]+\frac{36Lt}{n}\underset{i\leq n}{\max}K_{M}(i)+\frac{4L\sqrt{2t}}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}$ (32)
Let $\delta=\exp(-t)$, it follows from (31) and (32) that for all $n>N(c_{6},q)$, with at least probability $1-\delta$, we have

### v2

@@ 6 Discussion
In this paper, we proposed a framework of matching-based policy learning for identifying the optimal policy in observational studies by leveraging various matching estimators in the causal inference literature. Specifically, we developed a (bias-corrected) matching-based estimator of the advantage function and estimated the optimal policy by directly maximizing the resulting estimate over a class of policies. The proposed method possesses robustness in observational studies by circumventing the use of inverse probabilities. Meanwhile, covariate balancing is naturally attained by matching. Theoretically, we established a non-asymptotic bound for the regret of the learned policy, providing rigorous performance guarantee for MB-learning. Empirically, we demonstrated the competitive performance of the proposed method, particularly when extreme weights are present or when the sample size is not large.
There are several interesting directions in which this work can be further developed. First, MB-learning is intended for low-dimensional settings with binary treatment. In causal inference literature, Antonelli et al. (2018) proposed a doubly robust matching estimator to handle high-dimensional covariates, where variable selection techniques were used for matching with estimated propensity score and prognostic score. In addition, Wu et al. (2022) developed a matching estimator of average causal exposure-response function based on generalized propensity score. These estimators offer inspiration for adapting MB-learning to different scenarios. Second, we developed MB-learning with a fixed number of matches $M$, and our discussion in Section 3.2 demonstrated that a larger $M$ yields a lower regret bound. Lin et al. (2023) established an insightful result that the bias-corrected matching estimator (Abadie and Imbens, 2011) with a diverging $M$ exhibits doubly robustness and semiparametric efficiency under regularity conditions. MB-learning with a diverging $M$ may yield a doubly robust policy learning approach with an improved regret bound. The robustness of MB-learning may be partly attributed to covariate balance between the treatment groups, which is achieved through directly matching on covariates. It would be of considerable interest to integrate more sophisticated covariate balancing techniques for learning the optimal policy (e.g., Wong and Chan, 2018; Wang and Zubizarreta, 2020; Fan et al., 2022; Dai and Yan, 2024). Finally, extension of MB-learning to deal with multicategory or continuous policy is of future research interest (e.g., Li et al., 2025).
Supplementary Material
@@ Appendix A Technical Proofs
Before the proofs, we introduce some notations here. Let $(\Theta,d)$ be a semi-metric space, that is, the space $\Theta$ is equipped with a semi-metric $d$. For any $\epsilon>0$, let $N(\epsilon,\Theta,d)$ denote the covering number of $\Theta$, i.e., the minimal number of $d$-balls with radius $\epsilon$ needed to cover $\Theta$. We call a set of points is $\epsilon$-separated if the distance between every pair of points is strictly larger than $\epsilon$, and let $D(\epsilon,\Theta,d)$ denote the packing number of $\Theta$, i.e., the maximum number of points in $\Theta$ such that they are $\epsilon$-separated. The entropy numbers are the logarithms of covering and packing numbers respectively. These definitions can be found elsewhere (e.g., van der Vaart and Wellner, 2023). Throughout the proofs, we use $C$ to denote a generic positive constant, which may change from place to place.
@@ A.1 Proof of Lemma 1
Let $\mathbf{Y}=(Y_{1},\ldots,Y_{n})^{\top}$. Note that
@@
where the third equality follows from the fact that, if $M_{il}=1$, then $l\in\mathcal{J}_{M}(i)$ and thus $W_{i}=1-W_{l}$ by the definition of $\mathcal{J}_{M}(i)$. Combining these results, we have
@@ A.2 Proof of Lemma 2
Noting that $|K_{M}(\pi,i)|\leq\sum_{j=1}^{n}M_{ji}=K_{M}(i)$, the lemma follows directly from the Lemma 3 in Abadie and Imbens (2006). ∎
@@ A.3 Proof of Lemma 3
Let $\mathbf{\hat{Y}}(1)=\left(\hat{Y}_{1}(1),\ldots,\hat{Y}_{n}(1)\right)^{\top}$ and $\mathbf{\hat{Y}}(0)=\left(\hat{Y}_{1}(0),\ldots,\hat{Y}_{n}(0)\right)^{\top}$ denote the $n\times 1$ vectors of matching estimators for potential outcomes under each treatment. Then,
@@
Furthermore, let $\bm{\mu}\mathbf{(X\text{,}W)}=\left(\mu(X_{1},W_{1}),\ldots,\mu(X_{n},W_{n})\right)^{\top}$. It follows that
@@
Let $\bm{\phi}$ be the $n\times 1$ vector taking value in $\{-1,1\}^{n}$ with the $i$th element $12\pi(X_{i})-$ 1, and $\bm{\epsilon}$ be the $n\times 1$ vector with the $i$th element $\epsilon_{i}$ which is equal to $Y_{i}-\mu(X_{i},W_{i})$. Then, we obtain
@@
where the third equality follows from equation (7) and that $\mathbf{I_{n}}$ is the $n\times n$ identity matrix. Note that the elements of the two $n\times n$ matrices $\mathbf{G_{1}}-\mathbf{I_{n}}$ and $\mathbf{G_{0}}-\mathbf{I_{n}}$ are given by
@@
For each row $i$ such that $W_{i}=0$, we have $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}\bm{\mu}\mathbf{(X\text{,}1)}=\frac{1}{M}\sum_{j\in\mathcal{J}_{M}(i)}\mu(X_{j},1)-\mu(X_{i},1)$, where $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}$ is the $i$th row of $(\mathbf{G_{1}}-\mathbf{I_{n}})$; for each row $i$ such that $W_{i}=1$, we have $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}\bm{\mu}\mathbf{(X\text{,}1)}=0$. Hence,
@@
by similar arguments. Furthermore, it follows that
@@
After some algebra and similar arguments as the proof of Lemma 1, we have
@@ A.4 Proof of Lemma 4
Using the same argument in Abadie and Imbens (2006, Proof of Theorem 1, Page 258), we have $C^{2}n^{2/p}E\left[\frac{1}{n}\sum_{i=1}^{n}\|U_{i,M}\|^{2}\right]=O(1)$. Therefore, it holds that $B_{M}(\pi)=O_{p}\left(n^{-1/p}\right)$ by Markov’s inequality.
@@ A.5.1 Proof of Theorem 1 (i)
By Lemma 3, we rewrite
@@
We analyze each of the three terms respectively. For the first term, under the assumption that $\mu(x,w)$ is Lipschitz on a bounded set, then $\sup_{x,w}\mu(x,w)$ is bounded. By Weak Law of Large Numbers (van der Vaart, 1998, Proposition 2.16), $\bar{A}(\pi)-A(\pi)\stackrel{{\scriptstyle p}}{{\longrightarrow}}0$.
Let $\bar{\sigma}^{2}=\sup_{x,w}\sigma^{2}(x,w)$. It follows from the Lipschitz assumption that $\bar{\sigma}^{2}$ is bounded. Note that
@@
$\leq$ $E\left[\left(1+\frac{K_{M}(i)}{M}\right)^{2}\sigma^{2}(X_{i},W_{i})\right]$
$=$ $O(1),$
where the second equality follows from that for $i\neq j$, $E\left[\epsilon_{i}\epsilon_{j}|\{X_{i},W_{i}\}_{i=1}^{n}\right]=0$; the fourth inequality follows from triangle inequality and $|K_{M}(\pi,i)|\leq K_{M}(i)$; and the last equality follows from the boundedness of $\bar{\sigma}^{2}$ and Lemma 3 in Abadie and Imbens (2006). Hence, we obtain $E_{M}(\pi)=O_{p}\left(n^{-1/2}\right)=o_{p}(1)$ by Markov’s inequality. For the third term $B_{M}(\pi)$, it follows from Lemma 4 that $B_{M}(\pi)=O_{p}\left(n^{-1/p}\right)=o_{p}(1)$.
@@ A.5.2 Proof of Theorem 1 (ii)
Write
$\sqrt{n}\left(\hat{A}_{match}(\pi)-B_{M}(\pi)-A(\pi)\right)=\sqrt{n}\left(\bar{A}(\pi)-A(\pi)\right)+\sqrt{n}E_{M}(\pi).$
@@
For the second term, we consider the similar arguments in the proof of Abadie and Imbens (2006, Theorem 4). Let $E_{M,i}(\pi)=(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i}$. Conditional on $\{X_{i},W_{i}\}_{i=1}^{n}$, $K_{M}(\pi,i)$ is deterministic. Hence, the unit-level terms $\{E_{M,i}(\pi)\}^{n}_{i=1}$ are independently and not identically distributed with mean zero. We proceed to prove the asymptotic normality of $\sqrt{n}E_{M}(\pi)/\sqrt{V^{E}(\pi)}=\sum_{i=1}^{n}E_{M,i}(\pi)\sqrt{nV^{E}(\pi)}$ using Lindeberg-Feller Central Limit Theorem (Billingsley, 1995, Theorem 27.2). Specifically, given $\{X_{i},W_{i}\}_{i=1}^{n}$, we need to verify the Lindeberg condition:
@@ A.6 Proof of Theorem 2
We only need to verify that $\sqrt{n}\left(B_{M}(\pi)-\hat{B}_{M}(\pi)\right)=o_{p}(1)$. To complete the proof, we need to use the result of unit-level bias correction (Abadie and Imbens, 2011, Lemma A.2), which is given as follows:
@@ A.7 Proof of Theorem 3
Since $R(\hat{\pi})=\left(A(\pi^{*})-A(\hat{\pi})\right)/2$ by definition, it suffices to analyze $A(\pi^{*})-A(\hat{\pi})$ throughout the proof. Let $\Delta(\pi_{1},\pi_{2})=A(\pi_{1})-A(\pi_{2})$, $\bar{\Delta}(\pi_{1},\pi_{2})=\bar{A}(\pi_{1})-\bar{A}(\pi_{2})$, and $\hat{\Delta}(\pi_{1},\pi_{2})=\hat{A}^{bc}_{match}(\pi_{1})-\hat{A}^{bc}_{match}(\pi_{2})$. It follows from the triangle inequality that
@@
We will derive the non-asymptotic high probability bounds for the three terms in (A.7) in Step 1-3 separately, and then combine these bounds to complete the proof in Step 4.
@@
To derive this bound, we utilize Lemma 2 in Zhou et al. (2023). First, we rewrite $A(\pi)$ and $\bar{A}(\pi)$ in a vector-based form in order to apply their Lemma 2. Let
@@
We bound (I) using Hoeffding’s inequality (Vershynin, 2018, Theorem 2.2.6). Under the assumption that $|Y_{i}|\leq L$, we have $|\epsilon_{i}|=|Y_{i}-E[Y_{i}|X_{i},W_{i}]|\leq 2L$ by the triangle inequality. In addition, note that $E[(2W_{i}-1)\epsilon_{i}]=E\left[(2W_{i}-1)E\left[\epsilon_{i}\mid X_{i},W_{i}\right]\right]=0$. By Hoeffding’s inequality, for any $t>0$,
@@
Let $\delta=\exp\left(\frac{-t^{2}}{8L^{2}}\right)$. We obtain
@@
Let $\delta=\exp\left(\frac{-M^{2}t^{2}}{8L^{2}}\right)$. It follows that
@@
Let $\delta=\exp\left(-\frac{t^{2}}{64L^{2}}\right)$. Combining (20) and (21) we have
@@
It follows from the triangle inequality that
@@
Recall that in (26), we have showed that $G^{0}_{n}(\pi)$ is a sub-Gaussian process given $\{X_{i},W_{i},\epsilon_{i}\}^{n}_{i=1}$. Noting that $\|\tilde{\pi}_{k}-\tilde{\pi}_{k-1}\|_{n}\leq\|\tilde{\pi}_{k}-\pi\|_{n}+\|\tilde{\pi}_{k-1}-\pi\|_{n}\leq 3\tilde{D}2^{-k}$, we have
@@
Therefore, for some $c_{6}>0$, there exists $N(c_{6},q)$ such that for all $n>N(c_{6},q)$, we have $E\left[\max_{i\leq n}K_{M}(i)\right]\leq c_{6}n^{\frac{1}{q}}$. Combining (25), (30) and (31), for any $q>1$ and $n>N(c_{6},q)$, we have
@@
Let $b_{i}(\pi)=(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]$. Conditional on $\{X_{i},W_{i}\}_{i=1}^{n}$, the term
@@
is a sum of independent terms and $E\left[b_{i}(\pi)\epsilon_{i}\right]=0$. Note that
@@
$V=\underset{\pi\in\Pi}{\sup}\sum_{i=1}^{n}E_{\epsilon}\left[\frac{b^{2}_{i}(\pi)\epsilon^{2}_{i}}{4L^{2}\cdot\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}\right]\leq\underset{\pi\in\Pi}{\sup}\frac{\sum_{i=1}^{n}b^{2}_{i}(\pi)}{\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}\leq\frac{\sum_{i=1}^{n}K_{M}^{2}(i)}{\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}},$
@@
2($2\sqrt{\left(2V+16E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]\right)t}$
$\leq$ $2\sqrt{2Vt}+8\sqrt{E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]t}$
@@
Without loss of generality, we set $\alpha=1$. It follows that, conditional on $\{X_{i},W_{i}\}^{n}_{i=1}$, the following inequality holds with probability at least $1-\exp(-t)$:
@@
Since $\{b_{i}(\pi)\}_{i=1}^{n}$ are constants given $\{X_{i},W_{i}\}^{n}_{i=1}$, we replace the $E_{\epsilon}\left[\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]$ above by $E\left[\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]$. Taking expectation w.r.t. $\{X_{i},W_{i}\}^{n}_{i=1}$, for any $t>0$, with probability at least $1-\exp(-t)$, we have
$\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\leq 2E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]+\frac{36Lt}{n}\underset{i\leq n}{\max}K_{M}(i)+\frac{4L\sqrt{2t}}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}.$ (33)
Let $\delta=\exp(-t)$. It follows from (32) and (33) that for all $n>N(c_{6},q)$, with probability at least $1-\delta$, we have